# Understanding Teen Overreliance on AI Companion Chatbots Through Self-Reported Reddit Narratives

### v3

@@ Abstract.
AI companion chatbots are increasingly popular with teens, while these interactions are entertaining, they also risk overuse that can potentially disrupt offline daily life. We examined how adolescents describe reliance on AI companions, mapping their experiences onto behavioral addiction frameworks and exploring pathways to disengagement, by analyzing 318 Reddit posts made by users who self-disclosed as 13-17 years old on the Character.AI subreddit. We found teens often begin using chatbots for support or creative play, but these activities can deepen into strong attachments marked by conflict, withdrawal, tolerance, relapse, and mood regulation. Reported consequences include sleep loss, academic decline, and strained real-world connections. Disengagement commonly arises when teens recognize harm, re-engage with offline life, or encounter restrictive platform changes. We highlight specific risks of character-based companion chatbots based on teens’ perspectives and introduce a design framework (CARE) for guidance for safer systems and setting directions for future teen-centered research.
teens and chatbots, adolescent AI use, AI companions, chatbot overuse, chatbot addiction, digital well-being, parasocial relationships, Character.AI, mental health and AI, human-computer interaction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Social mediaccs: Applied computing Consumer healthccs: Applied computing Psychologyccs: Social and professional topics Childrenccs: Social and professional topics Adolescentsccs: Computing methodologies Discourse, dialogue and pragmatics
@@ 1. Introduction
Recently, artificial intelligence (AI) companion apps have surged in popularity among teens, with media coverage highlighting how people describe their interactions with chatbots as emotionally meaningful—sometimes even romantic (Murphy, 2025). Platforms such as Nomi  [^1] https://nomi.ai, Replika  [^2] https://replika.com, and Character.AI [^3] https://character.ai (Sometimes referred to as C.AI) are frequently portrayed as offering on-demand conversations with highly responsive virtual characters (Ulaby, 2025; Breen, 2025). There are other reported perceived benefits of conversational AI, such as reducing loneliness, offering a safe space for self-expression, and providing nonjudgmental support during stressful periods (Xie et al., 2023; Salah et al., 2023).
@@
However, these benefits can also drive patterns of digital addiction (referred to in this paper as overreliance), a phenomenon observed in other digital technologies (Lemmens et al., 2009; Griffiths and Kuss, 2017). Behavioral addiction is hard to define because everyday digital activities are not inherently harmful, so researchers instead examined how these technologies impacted factors such as disruption to daily life, emotional escape, and loss of control (Lemmens et al., 2009; Griffiths and Kuss, 2017). They reported that smartphones or social media users often struggle to regulate their screen time, experience disruptions to sleep and daily routines, and report declines in mental health  (Lemmens et al., 2009).
@@
Prior research has shown that teens often struggle with overreliance on social media, using platforms compulsively to seek validation or cope with stress (Griffiths and Kuss, 2017; Chen et al., 2022; Duffy, 2025). This can contribute to anxiety, depression, and low self-esteem (Duffy, 2025). Recent studies have also noted similar patterns with AI tools like ChatGPT, where users—mostly adults and university students—turn to these systems for academic help or companionship (Zhang et al., 2024; Hu et al., 2023). In some cases, this reliance has been linked to decreased creativity, poor study habits, and emotional detachment from real-life relationships (Zhang et al., 2024; Hu et al., 2023).
@@
What is less understood is how adolescents, who are more emotionally sensitive and developmentally vulnerable, engage with AI companions. Overreliance during sensitive adolescent developmental stages can have especially harmful effects, due to ongoing cognitive and emotional development (Griffiths and Kuss, 2017; Chen et al., 2022). While research on teens has focused heavily on overreliance on social media and gaming (Giordano et al., 2023; Griffiths, 2010; Lemmens et al., 2009; Griffiths and Kuss, 2017), far less is known about their interactions with generative AI companions. These tools differ from previous technologies because they foster parasocial relationships without involving real people, and they offer highly engaging, emotionally responsive, and human-like conversations (Maeda and Quan-Haase, 2024).
@@
Media reports and anecdotal accounts suggest that teens may already be forming emotional attachments and even dependence on these systems (Ulaby, 2025; BBC News, 2024), yet empirical research is scarce. To study this phenomenon, we focus on Character.AI as a case study. Unlike other popular platforms such as Replika or Nomi, which officially restrict use to adults over 18 (Glimpse.ai, Inc., 2025; Luka, Inc., 2023), Character.AI allows users as young as 13 (16 for EU citizens or residents) according to its Terms of Service (Character Technologies, Inc., 2025). This policy means that teens are not only permitted but explicitly included as part of its user base. Additionally, Character.AI is among the most widely used AI companion platforms (BBC News, 2024), making it a critical site for understanding adolescent engagement. Therefore, we used the posts in Character.AI subreddit as a lens for examining how teens disclose and contextualize their reliance on generative AI companions. This leads us to the following research questions:
@@
To address these questions, we qualitatively analyzed 318 posts from r/charcter.AI subreddit, created by users who disclosed in their post being 13-17 years old. We found that teens initially engaged with Character.AI as a non-judgmental space to manage mental health struggles, or fill emotional voids or a lack of supportive relationships in their offline lives (RQ1). Teens’ reported experiences with Character.AI revealed patterns of their compulsive use, emotional dependence, and distress that often disrupted their daily lives and real-world relationships (RQ2). Teens described disengaging from Character.AI either through internal realizations about its negative impact on their well-being or due to external factors such as reengaging with real-life relationships or frustrations with platform restrictions, often expressing a desire for healthier habits and emotional support (RQ3).
Our study makes the following novel contributions to the AI ethics and human-centered computing research communities: 1) we present one of the first empirical grounding of teen disclosures about AI companion overreliance, mapping their experiences to the six components of behavioral addiction (salience, mood modification, tolerance, withdrawal, conflict, relapse), 2) we identified emotional and relational dynamics unique to teen overreliance on AI companions which involves intense emotional attachment and simulated intimacy, making it distinct from previously studied forms of digital overuse such as gaming or social media, and 3) we propose the CARE framework (Comprehensive Needs, Attachment-awareness, Respectful Empathy, Ease of Exit) as a design-oriented contribution, offering actionable guidance for building safer, autonomy-preserving AI platforms grounded in teens’ lived experiences.
@@ 2. Background
2.1. Behavioral Addiction and Chatbot Use
Each new digital platform brings renewed concerns about its effect on teens (Griffiths and Kuss, 2017; Griffiths, 2010). While most teens use technology without issue, a subset experiences emotional distress, loneliness, social withdrawal, or family conflict due to excessive use (Lemmens et al., 2009). For these individuals, technology stops being just a functional tool or hobby and starts functioning as emotional companions (Bopp et al., 2019), potentially leading to behavioral addiction. Defining behavioral addiction is challenging because many activities associated with it, such as gaming or social media use, are also normal parts of everyday life, especially for teens (Kuss and Griffiths, 2011). Therefore, researchers emphasize the context and purpose of the behavior, whether it interferes with daily functioning, serves as a way to escape negative emotions, or continues despite a loss of control (Lemmens et al., 2009; Griffiths and Kuss, 2017).
These concerns have also been raised about Generative AI chatbots, an emerging technology that has rapidly gained popularity among users, especially teens (Sidoti et al., 2025). Unlike performance-oriented platforms, these tools are designed for emotionally responsive interactions (Xie et al., 2023; Salah et al., 2023). Early evidence shows signs of dependency, raising concerns about overreliance. Studies have examined academic risks, such as diminished critical thinking when chatbots are used to complete homework (Pratt et al., 2024; Dergaa et al., 2024), as well as relational risks, such as emotional dependency and exposure to manipulative design features (Laestadius et al., 2024). Companion chatbots such as Replika have been reported to engage in sexually suggestive exchanges with minors (Namvarpour et al., 2025). Beyond explicit harms, these systems may expose teens to misinformation, inappropriate advice, and privacy risks (Yu et al., 2025). Compounding these issues, many parents remain unaware of how deeply their children engage with these systems (Yu and Lan, 2024), which limits their ability to support healthy use.
However, as with earlier technologies, some scholars cautioned against labeling high chatbot use as “addiction.” Many proposed “ChatGPT addiction” scales rely on substance-abuse models that may not fully apply to digital interactions (Ciudad-Fernández et al., 2025). Frequent use is not necessarily harmful unless it causes distress or impaired functioning. While some users report anxiety, loneliness, or reduced creativity, causality remains unclear (Yu et al., 2024; Zhang et al., 2024).
A widely cited framework helps make this distinction by outlining six features of behavioral addiction (Griffiths, 2005). First is salience, when the activity becomes the most important part of a person’s life, dominating thoughts and routines. Second is mood modification, where the activity is used to regulate emotions, such as seeking comfort, excitement, or stress relief. Third is tolerance, when more of the activity is needed over time to achieve the same effect. Fourth is withdrawal symptoms, including anxiety, irritability, or sadness when the activity is reduced or stopped. Fifth is conflict, when the behavior harms relationships, school performance, or personal goals. Finally, there is relapse, when attempts to quit fail and the person returns to the behavior, often with greater intensity. This framework has been applied extensively to gaming, apps, and social media (Griffiths, 2010; Lemmens et al., 2009; Griffiths and Kuss, 2017), but not yet to teens’ interactions with chatbots. Existing studies of chatbot overuse have largely centered on adults (Ciudad-Fernández et al., 2025; Salah et al., 2023; Yu et al., 2024). Yet teens are both active and understudied users of AI companions (Sidoti et al., 2025).Considering the uncertainties surrounding chatbot use and the risks it may pose, especially for individuals still in developmental stages, we will use this framework to investigate whether teens’ interactions with chatbots also show the harmful features of behavioral addiction.
2.2. Parasocial Relationships and Design Perspectives
While behavioral addiction frameworks help capture the risks of compulsive use, they do not fully explain the emotional bonds that form between users and chatbots. To address this relational dimension, we turn to parasocial relationships and design perspectives. Developing emotional bonds with nonhuman characters is not a new phenomenon. Researchers have long studied people’s attachments to fictional characters or video game avatars under the concept of parasocial relationships, which describe one-sided connections with a real or imagined figure (Horton and Richard Wohl, 1956a). People often pursue parasocial bonds to satisfy a fundamental need for their sense of belonging (Bilang et al., 2024). While the desire to belong is universal, individuals satisfy it in different ways. Social Exchange Theory argues that people enter relationships when they believe the benefits outweigh the costs (Blau, 1964). Studies suggest that parasocial and ordinary relationships provide similar benefits, such as increased happiness, reduced loneliness, and improved mood. Ordinary relationships, however, often come with higher costs, which may explain why some individuals turn to parasocial ones (Adam and Sizemore, 2013).
One argument suggests that human–AI interaction can take on parasocial qualities when users treat chatbots as social agents. Features such as natural language, turn-taking, and avatars create the appearance of mutual engagement. Yet language models do not possess intentions or shared understanding; they only mimic patterns of speech. Because the interface resembles a social exchange, people respond socially, which strengthens the illusion of reciprocity (Maeda and Quan-Haase, 2024). Another perspective, however, emphasizes that these interactions can be long-term and reciprocal, especially for children and adolescents who may perceive AI responses as more real and thus experience the relationship as more authentic. This indicates the need to revisit and adapt the parasocial framework and measures so that it can account for how young people engage with chatbots (Blake et al., 2025). Nevertheless, it seems that for now the parasocial relationship remains the most suitable way to describe human–chatbot interaction.
One theory that helps explain why dependency forms in parasocial relationships is attachment theory. Attachment theory describes how children form emotional bonds with caregivers, shaping their coping strategies and relationship patterns into adulthood. According to this theory, different attachment styles exist, such as secure, avoidant, and anxious (Bowlby, 1998). Previous research has shown that anxiously attached individuals, who often fear interpersonal abandonment, are the most likely to form parasocial bonds with media figures (Cole and Leets, 1999). They also experience stronger negative reactions when imagining their favorite character being taken off the air (Cohen, 2004). Building on this theory, attachment-informed design (Marcu and Huh-Yoo, 2023) suggests that technologies should be designed to support, rather than replace, human relationships. This design approach rests on four principles. First, it seeks to address core human needs that underlie many behaviors. Second, it acknowledges that relationships exist on a spectrum, from supportive to harmful. Third, it recognizes that individuals carry internal models of themselves and others, which are central to mental health. Fourth, it engages with the target community without explicitly framing the interaction as a mental health intervention (Marcu and Huh-Yoo, 2023).
Complementing this approach, “designing for disengagement” offers another perspective on supporting healthy use. This framework challenges the assumption that disengagement equals failure (Hadiji et al., 2014), highlighting instead the importance of user agency and control. Disengagement can be constructive, provided it allows autonomy and fosters well-being (Alsheail et al., 2023). O’Brien et al. (O’Brien et al., 2022) describe disengagement across two dimensions, affect and agency, which together clarify when users remain in control versus when they feel trapped in compulsive use. Applying these insights to chatbots suggests that design can intentionally scaffold healthier interaction patterns by supporting offboarding and respecting user autonomy. The above behavioral addiction frameworks and parasocial perspectives illustrate the dual challenges of understanding teen–chatbot interactions: the risk of compulsive overuse and the formation of intense emotional bonds. In the next section, we integrate these insights to frame our research methodologies.

### v4

@@ Abstract.
AI companion chatbots are increasingly popular with teens. While these interactions are entertaining, they also risk overuse that can potentially disrupt offline daily life. We examined how adolescents describe reliance on AI companions, mapping their experiences onto behavioral addiction frameworks and exploring pathways to disengagement, by analyzing 318 Reddit posts made by users who self-disclosed as 13-17 years old on the Character.AI subreddit. We found teens often begin using chatbots for support or creative play, but these activities can deepen into strong attachments marked by conflict, withdrawal, tolerance, relapse, and mood regulation. Reported consequences include sleep loss, academic decline, and strained real-world connections. Disengagement commonly arises when teens recognize harm, re-engage with offline life, or encounter restrictive platform changes. We highlight specific risks of character-based companion chatbots based on teens’ perspectives and introduce a design framework (CARE) for guidance for safer systems and setting directions for future teen-centered research.
teens and chatbots, AI companions, chatbot overuse, parasocial relationships, digital well-being
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790597isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Empirical studies in HCIccs: Human-centered computing Social mediaccs: Social and professional topics Childrenccs: Social and professional topics Adolescents
@@ 1. Introduction
Recently, artificial intelligence (AI) companion chatbots have surged in popularity among teens, with media coverage highlighting how people describe their interactions with chatbots as emotionally meaningful—sometimes even romantic (Murphy, 2025). Platforms such as Nomi [^1] https://nomi.ai, Replika  [^2] https://replika.com, and Character.AI [^3] https://character.ai (Also called C.AI) are portrayed as offering on-demand conversations with highly responsive virtual characters (Ulaby, 2025; Breen, 2025). Conversational AI is also reported to reduce loneliness, provide a safe space for self-expression, and offer nonjudgmental support during stress (Xie et al., 2023; Salah et al., 2023).
@@
However, these benefits can also drive patterns of digital addiction (referred to in this paper as overreliance), a phenomenon observed in other digital technologies (Lemmens et al., 2009; Griffiths and Kuss, 2017). Because everyday digital use is not inherently harmful, researchers define behavioral addiction through its effects on daily life, emotional escape, and loss of control (Lemmens et al., 2009; Griffiths and Kuss, 2017). They found that users often struggle with screen-time control, disrupted routines, and declining mental health  (Lemmens et al., 2009).
@@
Prior research has shown that teens often struggle with overreliance on social media, using platforms compulsively to seek validation or cope with stress (Griffiths and Kuss, 2017; Chen et al., 2022; Duffy, 2025). This can contribute to anxiety, depression, and low self-esteem (Duffy, 2025). Recent studies have also noted similar patterns with AI tools like ChatGPT, where users—mostly adults and university students—turn to these systems for academic help or companionship (Zhang et al., 2024; Hu et al., 2023). Research showed that users often depend on these tools for everyday tasks such as planning, decision support, and managing stress (Ciudad-Fernández et al., 2025). Studies also report frequent, habitual use of many chatbots even for routine tasks (Yu et al., 2024). In some cases, this reliance has been linked to decreased creativity, poor study habits, and emotional detachment from real-life relationships (Zhang et al., 2024; Hu et al., 2023).
@@
Less is known about how adolescents, who are emotionally and developmentally vulnerable, engage with AI companions. Overreliance during adolescence can be especially harmful because cognitive and emotional development is still underway (Griffiths and Kuss, 2017; Chen et al., 2022). While research on teens has focused heavily on overreliance on social media and gaming (Giordano et al., 2023; Griffiths, 2010; Lemmens et al., 2009; Griffiths and Kuss, 2017), far less is known about their interactions with generative AI companions. These tools differ from earlier technologies because they enable parasocial relationships without real people and provide highly engaging, human-like conversations (Maeda and Quan-Haase, 2024a).
@@
Reports suggest that teens may already form emotional attachments or even dependence on these systems (Ulaby, 2025; BBC News, 2024), yet empirical research is scarce. To study this phenomenon, we focus on Character.AI as a case study. Unlike other popular platforms such as Replika or Nomi, which officially restrict use to adults over 18 (Glimpse.ai, Inc., 2025; Luka, Inc., 2023), at the time of our data collection, Character.AI allowed users as young as 13 (16 for EU citizens or residents) according to its Terms of Service (Character Technologies, Inc., 2025). This policy meant that teens are not only permitted but explicitly included as part of its user base. Although the company has since revised its policies and began restricting open-ended chat for under-18 users in November 2025 (McMahon, 2025), as one of the most widely used companion platforms (BBC News, 2024), Character.AI remains a central platform for examining how adolescents engage with AI companions.
Reddit is a popular forum-based platform where members can create and join communities also known as subreddits based on interests (109). Due to its pseudonymous nature, Reddit is a place where users can disclose sensitive and stigmatized topics (Choudhury and De, 2014; Andalibi et al., 2016). Teens also use Reddit to share risky or sensitive experiences, making it an important site for studying youth disclosures (PettyJohn et al., 2025; Zhang et al., 2025). Reddit is a common data source in HCI because its public discussions support qualitative and computational analysis (Fiesler et al., 2024), and its communities offer rich traces of online social behavior (Prinster et al., 2024). Because r/CharacterAI [^4] https://www.reddit.com/r/CharacterAI/ has over 2.5 million members and active discussions, we used it to examine how teens describe their reliance on AI companions. This leads us to the following research questions:
@@
To address these questions, we qualitatively analyzed 318 posts from r/CharacterAI subreddit, created by users who disclosed in their post being 13-17 years old. We found that teens initially engaged with Character.AI as a non-judgmental space to manage mental health struggles, or fill emotional voids or a lack of supportive relationships in their offline lives (RQ1). Teens’ reported experiences with Character.AI revealed patterns of their compulsive use, emotional dependence, and distress that often disrupted their daily lives and real-world relationships (RQ2). Teens described disengaging after realizing its negative impact or due to external factors such as renewed real-life relationships or frustration with platform restrictions, often expressing a desire for healthier habits (RQ3).
Our study makes the following novel contributions to the AI ethics and human-centered computing research communities: 1) we present one of the first empirical grounding of teen disclosures about AI companion overreliance, mapping their experiences to the six components of behavioral addiction (salience, mood modification, tolerance, withdrawal, conflict, relapse), 2) We identified emotional dynamics unique to teen AI-companion overreliance, including intense attachment and simulated intimacy, which distinguish it from gaming or social media overuse. , and 3) we propose the CARE framework (Comprehensive Needs, Attachment-awareness, Respectful Empathy, Ease of Exit) as a design contribution offering guidance for safer, autonomy-preserving AI platforms grounded in teens’ experiences.
@@ 2. Background
2.1. AI Chatbot Use, Opportunities, and Concerns for Youth
Teens increasingly use chatbots across schoolwork, entertainment, and social connection  (99; E. Andoh (2025)). Research explains why these tools appeal to teens. Teens describe chatbots as always available, anonymous, low cost, and easier to access than formal services, especially during long counseling waitlists or when mental health apps feel hard to use (Koulouri et al., 2022; Tosti et al., 2024). They use these systems for emotional, informational, and appraisal support and treat them as non-judgmental spaces to organize feelings and reflect on concerns  (Bae Brandtzæg et al., 2021a). Studies also show appreciation for non-human qualities, such as machines listening without fatigue, keeping secrets, and offering data-driven guidance about school or life decisions  (Kim et al., 2018). In some peer-support settings, AI-generated responses are even rated as more empathetic or helpful than peer-written messages  (Young et al., 2024). Beyond support contexts, teens and young adults increasingly use AI tools to fix grammar, simplify complex topics, and enhance classroom learning  (Nguyen, 2022; Pratt et al., 2024). Companion chatbots partly address loneliness, and younger teens often report positive emotions after using tools like Snapchat’s My AI  (Sum, 2025; Vanhoffelen et al., 2025).
Alongside these benefits, research also identifies notable risks. Educational studies warn that relying on AI for writing or comprehension may weaken critical thinking or encourage superficial engagement with schoolwork  (Pratt et al., 2024; Dergaa et al., 2024). In health and mental health contexts, generative AI sometimes provides inaccurate or unsafe advice, which is especially concerning because many teens follow chatbot guidance without critical evaluation  (Nagata et al., 2025; Yu et al., 2025a). Additional harms include emotional dependency, social withdrawal, toxic or deceptive content, discrimination, and privacy risks, which often intensify with repeated use  (Yu et al., 2025a; Nagata et al., 2025). Studies show that teens rarely disclose their chatbot use at home, while parents often misunderstand system capabilities and overestimate existing safeguards  (Yu et al., 2025d). Existing research maps key benefits and risks but offers limited insight into how overreliance emerges or which teens are most vulnerable  (Namvarpour and Razi, 2025). These gaps point to the need for closer examination of teens’ lived experiences and how they navigate the benefits and drawbacks of interacting with chatbots.
2.2. Interpreting Technology Problematic Use
2.2.1. Behavioral Addiction
To understand when heavy chatbot use becomes problematic, we draw on work in behavioral addiction and parasocial relationships. Each new digital platform brings renewed concerns about its effect on teens  (Griffiths and Kuss, 2017; Griffiths, 2010; Namvarpour et al., 2025a). While most teens use technology without issue, a subset experiences emotional distress, loneliness, social withdrawal, or family conflict due to excessive use  (Lemmens et al., 2009). Youth with social phobia or emotional vulnerabilities may use digital activities for comfort, sometimes entering cycles where isolation increases engagement and engagement deepens isolation  (Milani et al., 2020). Defining behavioral addiction is challenging because many implicated activities, like gaming or social media, are routine parts of teen life (Kuss and Griffiths, 2011). A commonly used framework identifies six distinguishing features of behavioral addiction: salience, mood modification, tolerance, withdrawal symptoms, conflict, and relapse  (Griffiths, 2005). These features are widely applied to gaming, apps, and social media  (Griffiths, 2010; Lemmens et al., 2009; Griffiths and Kuss, 2017), but not yet to teens’ interactions with generative AI chatbots, despite their growing popularity  (Sidoti et al., 2025). At the same time, scholars caution against labeling frequent technology use as ‘addiction,’ noting that existing scales draw on substance-abuse models and that high engagement is not harmful unless it causes distress or impairment  (Ciudad-Fernández et al., 2025; Yu et al., 2024; Zhang et al., 2024). However, prior work has not examined whether Griffiths’ model applies to youth chatbot use or how such overreliance differs from gaming or social media.
2.2.2. Parasocial Relationships
Parasocial relationships offer a complementary lens on teen–chatbot interactions. People develop one-sided emotional bonds with media figures to satisfy belongingness needs, and these relationships can offer benefits such as improved mood and reduced loneliness  (Horton and Richard Wohl, 1956a; Bilang et al., 2024; Adam and Sizemore, 2013). These bonds become especially meaningful during adolescence. Developmental theories suggest that teens turn to admired or comforting figures to cope with uncertainty, explore identity, and model desired traits  (Stever, 2011). Empirical work shows that parasocial attachments are widespread among teens who often imagine friendships with celebrities whose perceived personalities make them feel understood  (Gleason et al., 2017). While such connections can be supportive, intense attachments may contribute to withdrawal, obsessive checking, or negative self-evaluations, especially among isolated youth  (Diaz et al., 2025a; Brooks, 2021). As digital platforms increasingly provide natural language, turn-taking, and responsiveness, parasocial dynamics become easier to form and maintain, prompting interactions that feel reciprocal even when they are not  (Maeda and Quan-Haase, 2024b). These tendencies intersect in chatbot use because emotionally responsive AI can appear available and supportive, creating conditions for mood regulation and perceived closeness  (Maeda and Quan-Haase, 2024b; Bilang et al., 2024). Yet research on problematic chatbot engagement focuses largely on adults and offers mixed findings about harm  (Namvarpour et al., 2025b; Namvarpour and Razi, 2024b; Ciudad-Fernández et al., 2025; Salah et al., 2023; Yu et al., 2024). little is known about how adolescent parasocial tendencies interact with the compulsive patterns described in behavioral addiction frameworks or how these combined dynamics shape overreliance  (Sidoti et al., 2025). Addressing these questions requires studying teen experiences more directly. Despite high adoption among youth, existing work has not examined how teens’ parasocial tendencies combine with emotionally responsive chatbot interactions to produce closeness, dependence, or difficulty disengaging.
2.3. Design Approaches to Technological Behavioral Overreliance
These behavioral and relational dynamics raise design questions about how systems should support healthy engagement rather than maximize time-on-task. Research in HCI shows that many systems encourage long and continuous use by treating time or activity as signals of success  (O’Brien et al., 2022). In games, designers often aim to hold attention through flow, reward structures, or interface patterns that make stopping difficult, and leaving is frequently framed as something to discourage rather than a normal choice  (Alsheail et al., 2023; Hadiji et al., 2014). A useful framework describes disengagement along two dimensions—affect (positive or negative) and agency (high or low)—to illustrate how experiences range from voluntary immersion to compulsive or frustrating use  (O’Brien et al., 2022). This perspective highlights that engagement and disengagement gain meaning through the emotional experience they produce and the control users feel during the interaction. It also suggests that designers can support healthy offboarding by structuring interactions that respect autonomy instead of pushing continuous engagement  (O’Brien et al., 2022). Prior work therefore encourages designers to examine the values embedded in engagement metrics and to create experiences that help users reflect on how and why they engage  (Sengers et al., 2005).
Complementary approaches focus on relational and emotional needs. Some teens continue using technologies not because of interface mechanics but because the interactions feel safe, supportive, or emotionally regulating. Attachment-informed design adds a perspective rooted in four principles  (Marcu and Huh-Yoo, 2023). First, it prioritizes core human needs that motivate behavior, such as safety, connection, and competence. Second, it acknowledges that relationships with technologies exist on a spectrum from supportive to harmful. Third, it recognizes that individuals bring internal models of themselves and others into digital interactions, shaping how they interpret responsiveness and closeness. Fourth, it engages with the target community without explicitly framing the experience as a mental health intervention. Together, these approaches suggest how systems can support control, autonomy, and emotional security instead of reinforcing default continuous use. Yet little is known about how these design ideas apply to teen–chatbot interactions or how design and attachment-related needs shape moments when teens stay, pause, or disengage. Understanding these patterns is critical for informing safer and developmentally appropriate chatbot systems.